You are given the coordinates of the vertex (7,8) and of a point (-4,-9) on a parabola. Find the equation of the parabola
Found 2 solutions by nerdybill, oscargut:
Answer by nerdybill(7384)

You are given the coordinates of the vertex (7,8) and of a point (-4,-9) on a parabola. Find the equation of the parabola
.
Vertex form of parabola:
y= a(x-h)^2+k
plug provided info to find a:
8= a(7-(-4))^2+(-9)
8= a(7+4)^2+(-9)
8= a(11)^2+(-9)
17= a(11)^2
17= a(121)
17/121 = a
.
Equation then is:
y= (17/121)(x-(-4))^2+(-9)
y= (17/121)(x+4)^2-9



Answer by oscargut(2103)

Vertex is (7,8)
then equation is:
y = a(x-7)^2+8
(-4,-9) is in the parabola then:
-9 = a(-4-7)^2+8
-9 = 121a+8
121a = -17
Answer: y = (-17/121)(x-7)^2+8
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com